Assume buying price from the farmer is $0.50 per kg, the weighted average wholesale selling price is $0.90 per kg and the weighted average retail price is $1.17 per kg.caoculate the total Market margin


1
Expert's answer
2021-08-03T11:51:48-0400

Total margin= total revenue-total expenses/total revenue

Total Market Margin = 1.17- 0.5=0.67/1.17

= 0.572


0.90-0.05= 0.4/0.9

=0.4444